OpenStudy (anonymous):

y(n)-5y(n-1)+6y(n-2)=0

OpenStudy (anonymous):

we have to solve for y(n)

OpenStudy (anonymous):

so I took fourier transform \[y(e^{jw})-5y(e^{jw})e^{-jw}+6y(e^{jw})e^{-2jw}=0\] \[y(e^{jw})(1-5e^{-jw}+6e^{-2jw})=0\]

OpenStudy (anonymous):

\[y(e^{jw})(1-5e^{-jw}+6e^{-2jw})=0\] let s=e^jw \[y(s)(1-5/s +6/s^2)=0\]

OpenStudy (anonymous):

\[(1-5/s +6/s^2)=0\] \[(1-3/s)(1-2/s)=0\]

OpenStudy (anonymous):

\[(1-3s^{-1})(1-2s^{-1})=0\]

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

s=-3,-2

OpenStudy (phi):

I think the roots are +3 and +2 (I would have solved s^2 -5s +6=0 to get (s-2)(s-3)=0

OpenStudy (anonymous):

yes, s=3 and s=2

OpenStudy (anonymous):

then how would we go back to fourier form?

OpenStudy (phi):

this reminds me of laplace transforms, but I have not seen it done using fourier

OpenStudy (anonymous):

z transform since it is discreet

OpenStudy (phi):

I would have to read up on it, because I don't know this off the top of my head.

OpenStudy (anonymous):

ok, thanks

OpenStudy (phi):

btw, what kind of problem is this?

OpenStudy (anonymous):

this is from discrete signal processing

OpenStudy (phi):

The best I can come up with is that you are solving a homogenous equation for the non-forced system response. The solution is a linear combination of exponentials. you found \[ e^{jw_0} = 2 \text{ and } e^{jw_1} =3 \] the solution would be \[y= c_0 e^{jw_0n} +c_1 e^{jw_1n} \] which becomes \[y= c_0 2^{n} +c_1 3^{n} \]